OpenStudy (anonymous):

how would you find the regular form of a line with a slope of 2, that passes through the point (0,-2)? Not remembering basic math..

OpenStudy (anonymous):

use the point slope formula: \[y-y_1=m(x-x_2)\] where x\[x_1, y_1\] are (0,-2) and m is the slope = 2

OpenStudy (anonymous):

therefore: \[y-(-2)=2(x-0)\]

OpenStudy (anonymous):

multiply through and move everything but the y to the right side of the equation: y+2=2x-2(0) y=2x-0-2 y=2x-2

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

No Probrrrem :)